OpenStudy (anonymous):

Write the equation of the line given the slope and a point on the line. Type your answer in the box without skipping any spaces. m = -4, point (2, -5)

OpenStudy (anonymous):

You would use the point-slope form of an equation:\[y - y _{1} = m(x - x _{1})\]where you are given point (x1, y1) and your slope "m".

OpenStudy (anonymous):

(y- - 5) = -4(2-x1) would that be correct?

OpenStudy (anonymous):

i mean (y--5)=-4(x-2) i mean

OpenStudy (anonymous):

& that would be y=-4x+3?

OpenStudy (anonymous):

yes. Good job!

OpenStudy (anonymous):

thanks!